# Universidade de Lisboa, Av. Rovisco Pais 1, 1049-001 Lisboa, Portugal

### v3

@@ 1 Introduction
Despite the large success of the Standard Model (SM) of particle physics in providing experimental predictions, leading to the discovery of a new scalar particle resembling the predicted Higgs boson ATLAS:2012yve; CMS:2012qbp, there is a general consensus that there must be Physics Beyond the Standard Model (BSM). The critical open problems consist of: the need for new sources of CP-violation, a necessary ingredient for a successful explanation of the baryon asymmetry of the universe Sakharov:1967dj; an explanation for the existence and nature of Dark Matter (DM), which comprises of order 85 % of the matter content of the Universe Planck:2013oqw; and the origin of the observed tiny neutrino masses. Motivated by the lack of a fundamental reason why the scalar sector should be limited to a single Higgs doublet, Higgs-sector extensions are required for many of the viable explanations to these problems.
@@
In this work, we consider a two Higgs doublet model (2HDM) Branco:2011iw with a $\mathbbm{Z}_{2}$ symmetry imposed in order to forbid Higgs-mediated flavor changing neutral couplings (FCNCs) at the tree-level, with fermions of a given electric charge coupling to only one Higgs doublet. Of the four possible choices, the type II 2HDM is the most studied, since it corresponds to the structure present in supersymmetric models - see, for example, refs. Djouadi:2005gj; Dreiner:2023yus. We, however, consider the general type II 2HDM and aim to extend the scalar sector with additional singlets as viable candidates for particle Dark Matter. When protected by a new global symmetry that remains unbroken in the vacuum, the singlets can meet the requirements of heavy, stable, electrically neutral particles of non-baryonic nature Dimopoulos:1989hk; McDermott:2010pa. We will consider a scenario where the DM particle are produced via the freeze-out mechanism Bertone:2004pz; Feng:2010gw. A particle candidate with a mass similar to the scale of electroweak symmetry breaking and an interaction cross section with the SM particles of the order of the weak force processes can meet the observed relic abundance, and fits the class of Weakly Interacting Massive Particles (WIMPs). The model is then subjected to all theoretical, collider, astrophysical, and cosmological constraints to obtain the allowed parameter space region.
@@
The 2HDM extended by an additional real singlet with a $\mathbbm{Z}_{2}$ parity symmetry may contain a viable Dark Matter candidate when the singlet does not acquire a vacuum expectation value (vev), see e.g. refs. Grzadkowski:2009iz; Logan:2010nw; Boucenna:2011hy; He:2011gc; Bai:2012nv; Drozd:2014yla; Campbell:2015fra. It has been dubbed the Next-to-Two Higgs Doublet model (N2HDM) in Chen:2013jvg; Muhlleitner:2016mzt; Ferreira:2019iqb; Engeln:2020fld; Glaus:2022rdc. In 2022, there was already a tension in the N2HDM between the relic density measured by PLANCK Planck:2018vyg and exclusion bounds from DM direct detection experiments, as shown in figure 3 of Glaus:2022rdc. Since then, the situation has worsened, given the great improvement in the scattering bounds LZ:2024zvo.
@@
A possible alternative is to consider a Type II 2HDM augmented with a complex scalar singlet (2HDMS), either without a DM candidate Baum:2018zhf; Heinemeyer:2021msz, or with a DM candidate enabled through a suitable symmetry and vacuum Dutta:2022qeq; Dutta:2023cig; Lahiri:2024rya. Such studies have concentrated thus far on benchmarks rather than dedicated multi-variable scans.
@@
Extensions of the scalar sector of the 2HDM through the addition of more than one singlet offer promising frameworks that could also address the Dark Matter puzzle  [^1] The interesting possibility of adding multiple singlets to the (one Higgs) SM, has also been considered. For a recent analysis see, for example, Ref. Goncalves:2025snm.. Within this context, we consider a two real scalar inert singlet extension of the two Higgs doublet model (2HDM2S). It includes all theoretical derivations relevant for unitarity, boundedness from below, and vacuum stability. It also includes a full parameter scan simulation of the model, performed in three ways: i) with a traditional scan of the full parameter space; ii) starting from scans close to the alignment limit; and, iii) using a Machine Learning optimization technique, evoking Evolutionary Strategies.
@@ 2 The Scalar Potential of the 2HDM2S
Our goal is to analyze an extension to the SM, with two doublets and two real scalars, in which the scalars do not acquire a Vacuum Expectation Value (vev) after Spontaneous Symmetry Breaking (SSB). In order to avoid dangerous flavor-changing neutral couplings (FCNC) at the tree level, a $\mathbbm{Z}_{2}$ symmetry softly broken by the $m_{12}^{2}$ term, to introduce a decoupling limit Gunion:2002zf, is imposed on the Lagrangian. In addition, we assume CP conservation in the theory; therefore, all the coefficients in the potential are taken to be real.
@@
where we have defined Faro:2019vcd,
@@
where $\tan\beta=\frac{v_{2}}{v_{1}}$, $s_{\beta}=\sin(\beta)$, and $c_{\beta}=\cos(\beta)$. With this definition we may obtain the physical-mass eigenstates Das:2019yad by
@@ 3.1 Stationarity Conditions
The most general constant field configuration for the vacuum is, using the notation from Muhlleitner:2016mzt; Ferreira:2019iqb,
@@ 3.2 Formalism and Vacua
To study the interplay of multiple vacua, it is useful to introduce a bilinear formalism, similar to the one used in Lahiri:2024rya; Ferreira:2019iqb for the 2HDMS and N2HDM. The relevant bilinears for this work are the following.
@@ 4 Sufficient BFB conditions
We now will use a parameterization inspired by Ref. Faro:2019vcd,
@@
The problem is that we have to find the conditions for a matrix of order four to be definite positive. For the case of matrices of order three there is a simple result. A $3\times 3$ symmetric matrix is positive definite if and only if the following conditions, known as copositivity conditions Klimenko:1984qx; Kannike:2012pe, are satisfied:
@@
We will come back to these expressions after we find a quadratic form that bounds the potential from below. As the odd part, $V_{O}$, cannot be written as a quadratic form, one has to find sufficient although not necessary conditions for the BFB Boto:2022uwv. For this, we bound each part of the potential. We have
@@
is positive definite. The criteria for verifying the copositivity of the $4\times 4$ matrix $\bar{A}$ are given in Ref. PING1993109. We have implemented this long algorithm (it has 49 steps) and verified, using the CERN library Minuit james:1975dr, that all points that passed the above criteria were indeed BFB.
@@ 5.3 $v_{1}=0,v_{2}=0,v_{P}=0$ and $v_{S}\not=0$
Then, if $v_{S}^{2}>0$, the potential at the the minimum is given by,
@@ 5.4 $v_{1}=0,v_{S}=0$ and $v_{2}\not=0,v_{P}\not=0$
Then, if $v_{P}^{2}>0$ and $v_{2}^{2}>0$, the potential at the the minimum is given by,
@@ 5.5 $v_{1}=0,v_{P}=0$ and $v_{2}\not=0,v_{S}\not=0$
Then, if $v_{S}^{2}>0$ and $v_{2}^{2}>0$, the potential at the the minimum is given by,
@@ 5.6 $v_{2}=0,v_{S}=0$ and $v_{1}\not=0,v_{P}\not=0$
Then, if $v_{P}^{2}>0$ and $v_{1}^{2}>0$, the potential at the the minimum is given by,
@@ 5.7 $v_{2}=0,v_{P}=0$ and $v_{1}\not=0,v_{S}\not=0$
Then, if $v_{S}^{2}>0$ and $v_{1}^{2}>0$, the potential at the the minimum is given by,
@@ 5.8 $v_{2}=0,v_{P}=0,v_{S}=0$ and $v_{1}\not=0$
Then, if $v_{1}^{2}>0$, the potential at the the minimum is given by,
@@ 5.9 $v_{1}=0,v_{P}=0,v_{S}=0$ and $v_{2}\not=0$
Then, if $v_{2}^{2}>0$, the potential at the the minimum is given by,
@@ 5.10 Other minima
So, if our minimum, in eq. (80), is above one of the cases in eqs. (82),(84), (86), (88), (90), (92), (94),(96) then the point in parameter space must be discarded. However, the list above does not contain all the possible minima. We have verified with a minimization procedure using Minuit james:1975dr that there are other two cases. These do not have an analytical solution (system of coupling cubic non-linear equations), but we have found a way of solving them numerically. We will discuss these in the next two subsections.
@@ 6 Perturbative Unitarity
For the perturbative unitarity, we follow section 4 of Ref. Bento:2017eti. To be self-contained we reproduce here their table 3 for our case. For convenience we define
@@
and we get the results in table 1. Notice that in comparison with Table 3 of Ref. Bento:2017eti, we have
@@
Applying the procedure described in Ref. Bento:2017eti we get the following results with the notation $M_{\mathcal{Y}}^{\mathcal{Q}}$.
@@
The eigenvalues separate in six that can be easily evaluated, but the other five are solutions of a fifth order polynomial. In this case it is easier to use the methods of the minors as explained in Ref. Bento:2022vsb for the whole matrix.
@@ 7 The precision observables $S,T$ and $U$
To obtain these observables one follows the setup of Ref. Grimus:2007if. For this one has to obtain the matrices $U$ and $V$ in their notation.
Following Ref. Grimus:2007if, and using the definitions given in section 2 we define
@@
We have now everything to define the matrices $U_{2\times 2}$, $V_{2\times 6}$ and $R_{2\times 6}.$ [^3] As can be seen from the expressions for $S,T,U$ in Ref. Grimus:2007if, the matrix $R$ is not necessary. We get,
@@
Using these matrices, we have programmed the expressions in Ref. Grimus:2007if. The only important point to remember is the order of the fields in eq. (126).
@@ 8 Experimental Constraints
We have discussed the conditions for boundedness from below, a global minimum and perturbative unitarity. We complete the theoretical constraints by requiring perturbativity of the Type-II Yukawa couplings, by setting them to be $|y_{i}|<\sqrt{4\pi}$, where $i=t,b,\tau$. The bounds from experimental collider searches follow: the oblique parameters STU described in Section 7 are to be compared with the global electroweak fit in Baak:2014ora; the coupling-strength modifiers within $3\sigma$ of the LHC data ATLAS:2019nkf; the LHC signal strengths of the $125\textrm{GeV}$ Higgs, for the combinations of production cross sections and branching ratios, to have a $2\sigma$ agreement with the most recent ATLAS results ATLAS:2022vkf. For LHC searches for new particles, we use the software package HiggsTools-1.1.3 Bahl:2022igd, which includes the latest data from the ATLAS and CMS experiments at CERN, and the latest LHC bound on the invisible branching ratio of the Higgs, BR($h\rightarrow\text{inv}$) $\leq$ 0.107 at $95\%$ C.L. ATLAS:2023tkt.
For the DM observables, we implemented our model in micrOMEGAs-6.2.3 Alguero:2023zol to numerically calculate the relic density, scattering amplitudes, annihilation cross section and decay rates. We note that the $\mathbbm{Z}_{2}$ symmetry used in this work does not separate the DM particle into two non-communicating sectors. One could think that a second DM particle could be invoked by imposing a kinematic limit on the mass of the second DM candidate. However, in this case, such putative DM candidate could decay into the (true) lightest DM candidate and a SM pair, in particular through the terms proportional to $\lambda_{15}$ and $\lambda_{16}$ in Eq. (3). Such decays are included using the micrOMEGAs routine darkOmegaN but not under the darkOmega2 routine Alguero:2023zol, which comes activated by default in the usual micrOMEGAs distribution. We compare with the Planck experiment Planck:2018vyg at $3\sigma$,
@@
For the DM-nucleon scattering, we follow the method described in Ref. Belanger:2014bga of computing the normalized cross section of DM on a point-like Xenon nucleus
@@
with $\mu_{\text{DM}}$ the reduced mass of the DM candidate and $f_{p}$, $f_{n}$ the amplitudes for protons and neutrons. We compare with the most recent LZ release in 2024 LZ:2024zvo.
The annihilation cross section calculated is to be compared with reconstructions based on indirect searches. We follow the method described in Belanger:2021lwd and compare the dominating channel with the respective experimental exclusion line. We find that for our model and mass ranges, the annihilation occurs either into $VV$, summing only the $WW$ and $ZZ$ final states, dubbed $\langle\sigma v\rangle_{VV}$, or into $b\bar{b}$  [^4] There are currently no published equivalent exclusion lines for decays going mainly into $c\bar{c}$ nor $h_{1}h_{1}$.. For each point in parameter space, the procedure is to check the annihilation channel and apply the respective exclusion bounds in Fig. 1. We have confirmed that for the mass region studied, this constraint does not exclude a significant amount of the parameter space when applied after demanding the correct relic density.
Figure 1: Combined relevant limits from indirect searches on the total $\langle\sigma v\rangle$ as a function of the mass of the DM candidate $m_{DM}$. The lines coming from Fermi-LAT Fermi-LAT:2015att and H.E.S.S. HESS:2022ygk assume a Navarro-Frenk-White (NFW) DM density profile and the AMS-02 AMS:2016oqu lines correspond to the conservative approach derived in Ref. Reinert:2017aga.

### v4

@@ 1 Introduction
Despite the large success of the Standard Model (SM) of particle physics in providing experimental predictions, leading to the discovery of a new scalar particle resembling the predicted Higgs boson Aad and others (2012); Chatrchyan and others (2012), there is a general consensus that there must be Physics Beyond the Standard Model (BSM). The critical open problems consist of: the need for new sources of CP-violation, a necessary ingredient for a successful explanation of the baryon asymmetry of the universe Sakharov (1967); an explanation for the existence and nature of Dark Matter (DM), which comprises of order 85 % of the matter content of the Universe Ade and others (2014); and the origin of the observed tiny neutrino masses. Motivated by the lack of a fundamental reason why the scalar sector should be limited to a single Higgs doublet, Higgs-sector extensions are required for many of the viable explanations to these problems.
@@
In this work, we consider a two Higgs doublet model (2HDM) Branco et al. (2012) with a $\mathbbm{Z}_{2}$ symmetry imposed in order to forbid Higgs-mediated flavor changing neutral couplings (FCNCs) at the tree-level, with fermions of a given electric charge coupling to only one Higgs doublet. Of the four possible choices, the type II 2HDM is the most studied, since it corresponds to the structure present in supersymmetric models - see, for example, refs. Djouadi (2008); Dreiner et al. (2023). We, however, consider the general type II 2HDM and aim to extend the scalar sector with additional singlets as viable candidates for particle Dark Matter. When protected by a new global symmetry that remains unbroken in the vacuum, the singlets can meet the requirements of heavy, stable, electrically neutral particles of non-baryonic nature Dimopoulos et al. (1990); McDermott et al. (2011). We will consider a scenario where the DM particle are produced via the freeze-out mechanism Bertone et al. (2005); Feng (2010). A particle candidate with a mass similar to the scale of electroweak symmetry breaking and an interaction cross section with the SM particles of the order of the weak force processes can meet the observed relic abundance, and fits the class of Weakly Interacting Massive Particles (WIMPs). The model is then subjected to all theoretical, collider, astrophysical, and cosmological constraints to obtain the allowed parameter space region.
@@
The 2HDM extended by an additional real singlet with a $\mathbbm{Z}_{2}$ parity symmetry may contain a viable Dark Matter candidate when the singlet does not acquire a vacuum expectation value (vev), see e.g. refs. Grzadkowski and Osland (2010); Logan (2011); Boucenna and Profumo (2011); He et al. (2012); Bai et al. (2013); Drozd et al. (2014); Campbell et al. (2015). It has been dubbed the Next-to-Two Higgs Doublet model (N2HDM) in Chen et al. (2014); Muhlleitner et al. (2017); Ferreira et al. (2019); Engeln et al. (2020); Glaus et al. (2022). In 2022, there was already a tension in the N2HDM between the relic density measured by PLANCK Aghanim and others (2020) and exclusion bounds from DM direct detection experiments, as shown in figure 3 of Glaus et al. (2022). Since then, the situation has worsened, given the great improvement in the scattering bounds Aalbers and others (2024a).
@@
A possible alternative is to consider a Type II 2HDM augmented with a complex scalar singlet (2HDMS), either without a DM candidate Baum and Shah (2018); Heinemeyer et al. (2022), or with a DM candidate enabled through a suitable symmetry and vacuum Dutta et al. (2025, 2024); Lahiri and Moortgat-Pick (2024). Such studies have concentrated thus far on benchmarks rather than dedicated multi-variable scans.
@@
Extensions of the scalar sector of the 2HDM through the addition of more than one singlet offer promising frameworks that could also address the Dark Matter puzzle  [^1] The interesting possibility of adding multiple singlets to the (one Higgs) SM, has also been considered. For a recent analysis see, for example, Ref. Gonçalves et al. (2025).. Within this context, we consider a two real scalar inert singlet extension of the two Higgs doublet model (2HDM2S). It includes all theoretical derivations relevant for unitarity, boundedness from below, and vacuum stability. It also includes a full parameter scan simulation of the model, performed in three ways: i) with a traditional scan of the full parameter space; ii) starting from scans close to the alignment limit; and, iii) using a Machine Learning optimization technique, evoking Evolutionary Strategies.
@@ 2 The Scalar Potential of the 2HDM2S
Our goal is to analyze an extension to the SM, with two doublets and two real scalars, in which the scalars do not acquire a Vacuum Expectation Value (vev) after Spontaneous Symmetry Breaking (SSB). In order to avoid dangerous flavor-changing neutral couplings (FCNC) at the tree level, a $\mathbbm{Z}_{2}$ symmetry softly broken by the $m_{12}^{2}$ term, to introduce a decoupling limit Gunion and Haber (2003), is imposed on the Lagrangian. In addition, we assume CP conservation in the theory; therefore, all the coefficients in the potential are taken to be real.
@@
where we have defined Faro and Ivanov (2019),
@@
where $\tan\beta=\frac{v_{2}}{v_{1}}$, $s_{\beta}=\sin(\beta)$, and $c_{\beta}=\cos(\beta)$. With this definition we may obtain the physical-mass eigenstates Das and Saha (2019) by
@@ 3.1 Stationarity Conditions
The most general constant field configuration for the vacuum is, using the notation from Muhlleitner et al. (2017); Ferreira et al. (2019),
@@ 3.2 Formalism and Vacua
To study the interplay of multiple vacua, it is useful to introduce a bilinear formalism, similar to the one used in Lahiri and Moortgat-Pick (2024); Ferreira et al. (2019) for the 2HDMS and N2HDM. The relevant bilinears for this work are the following.
@@ 4 Sufficient BFB conditions
We now will use a parameterization inspired by Ref. Faro and Ivanov (2019),
@@
The problem is that we have to find the conditions for a matrix of order four to be definite positive. For the case of matrices of order three there is a simple result. A $3\times 3$ symmetric matrix is positive definite if and only if the following conditions, known as copositivity conditions Klimenko (1985); Kannike (2012), are satisfied:
@@
We will come back to these expressions after we find a quadratic form that bounds the potential from below. As the odd part, $V_{O}$, cannot be written as a quadratic form, one has to find sufficient although not necessary conditions for the BFB Boto et al. (2022). For this, we bound each part of the potential. We have
@@
is positive definite. The criteria for verifying the copositivity of the $4\times 4$ matrix $\bar{A}$ are given in Ref. Ping and Yu (1993). We have implemented this long algorithm (it has 49 steps) and verified, using the CERN library Minuit James and Roos (1975), that all points that passed the above criteria were indeed BFB.
@@ 5.3 $v_{1}=0,v_{2}=0,v_{P}=0$ and $v_{S}\not=0$
Then, if $v_{S}^{2}>0$, the potential at the minimum is given by,
@@ 5.4 $v_{1}=0,v_{S}=0$ and $v_{2}\not=0,v_{P}\not=0$
Then, if $v_{P}^{2}>0$ and $v_{2}^{2}>0$, the potential at the minimum is given by,
@@ 5.5 $v_{1}=0,v_{P}=0$ and $v_{2}\not=0,v_{S}\not=0$
Then, if $v_{S}^{2}>0$ and $v_{2}^{2}>0$, the potential at the minimum is given by,
@@ 5.6 $v_{2}=0,v_{S}=0$ and $v_{1}\not=0,v_{P}\not=0$
Then, if $v_{P}^{2}>0$ and $v_{1}^{2}>0$, the potential at the minimum is given by,
@@ 5.7 $v_{2}=0,v_{P}=0$ and $v_{1}\not=0,v_{S}\not=0$
Then, if $v_{S}^{2}>0$ and $v_{1}^{2}>0$, the potential at the minimum is given by,
@@ 5.8 $v_{2}=0,v_{P}=0,v_{S}=0$ and $v_{1}\not=0$
Then, if $v_{1}^{2}>0$, the potential at the minimum is given by,
@@ 5.9 $v_{1}=0,v_{P}=0,v_{S}=0$ and $v_{2}\not=0$
Then, if $v_{2}^{2}>0$, the potential at the minimum is given by,
@@ 5.10 Other minima
So, if our minimum, in eq. (80), is above one of the cases in eqs. (82),(84), (86), (88), (90), (92), (94),(96) then the point in parameter space must be discarded. However, the list above does not contain all the possible minima. We have verified with a minimization procedure using Minuit James and Roos (1975) that there are other two cases. These do not have an analytical solution (system of coupling cubic non-linear equations), but we have found a way of solving them numerically. We will discuss these in the next two subsections.
@@ 6 Perturbative Unitarity
For the perturbative unitarity, we follow section 4 of Ref. Bento et al. (2017). To be self-contained we reproduce here their table 3 for our case. For convenience we define
@@
and we get the results in table 1. Notice that in comparison with Table 3 of Ref. Bento et al. (2017), we have
@@
Applying the procedure described in Ref. Bento et al. (2017) we get the following results with the notation $M_{\mathcal{Y}}^{\mathcal{Q}}$.
@@
The eigenvalues separate in six that can be easily evaluated, but the other five are solutions of a fifth order polynomial. In this case it is easier to use the methods of the minors as explained in Ref. Bento et al. (2022) for the whole matrix.
@@ 7 The precision observables $S,T$ and $U$
To obtain these observables one follows the setup of Ref. Grimus et al. (2008). For this one has to obtain the matrices $U$ and $V$ in their notation.
Following Ref. Grimus et al. (2008), and using the definitions given in section 2 we define
@@
We have now everything to define the matrices $U_{2\times 2}$, $V_{2\times 6}$ and $R_{2\times 6}.$ [^3] As can be seen from the expressions for $S,T,U$ in Ref. Grimus et al. (2008), the matrix $R$ is not necessary. We get,
@@
Using these matrices, we have programmed the expressions in Ref. Grimus et al. (2008). The only important point to remember is the order of the fields in eq. (126).
@@ 8 Experimental Constraints
We have discussed the conditions for boundedness from below, a global minimum and perturbative unitarity. We complete the theoretical constraints by requiring perturbativity of the Type-II Yukawa couplings, by setting them to be $|y_{i}|<\sqrt{4\pi}$, where $i=t,b,\tau$. The bounds from experimental collider searches follow: the oblique parameters STU described in Section 7 are to be compared with the global electroweak fit in Baak et al. (2014); the coupling-strength modifiers within $3\sigma$ of the LHC data Aad and others (2020); the LHC signal strengths of the $125\textrm{GeV}$ Higgs, for the combinations of production cross sections and branching ratios, to have a $2\sigma$ agreement with the most recent ATLAS results Aad and others (2022). For LHC searches for new particles, we use the software package HiggsTools-1.1.3 Bahl et al. (2023), which includes the latest data from the ATLAS and CMS experiments at CERN, and the latest LHC bound on the invisible branching ratio of the Higgs, BR($h\rightarrow\text{inv}$) $\leq$ 0.107 at $95\%$ C.L. Aad and others (2024).
For the DM observables, we implemented our model in micrOMEGAs-6.2.3 Alguero et al. (2024) to numerically calculate the relic density, scattering amplitudes, annihilation cross section and decay rates. We note that the $\mathbbm{Z}_{2}$ symmetry used in this work does not separate the DM particle into two non-communicating sectors. One could think that a second DM particle could be invoked by imposing a kinematic limit on the mass of the second DM candidate. However, in this case, such putative DM candidate could decay into the (true) lightest DM candidate and a SM pair, in particular through the terms proportional to $\lambda_{15}$ and $\lambda_{16}$ in Eq. (3). Such decays are included using the micrOMEGAs routine darkOmegaN but not under the darkOmega2 routine Alguero et al. (2024), which comes activated by default in the usual micrOMEGAs distribution. We compare with the Planck experiment Aghanim and others (2020) at $3\sigma$,
@@
For the DM-nucleon scattering, we follow the method described in Ref. Bélanger et al. (2014) of computing the normalized cross section of DM on a point-like Xenon nucleus
@@
with $\mu_{\text{DM}}$ the reduced mass of the DM candidate and $f_{p}$, $f_{n}$ the amplitudes for protons and neutrons. We compare with the most recent LZ release in 2024 Aalbers and others (2024a).
The annihilation cross section calculated is to be compared with reconstructions based on indirect searches. We follow the method described in Belanger et al. (2022) and compare the dominating channel with the respective experimental exclusion line. We find that for our model and mass ranges, the annihilation occurs either into $VV$, summing only the $WW$ and $ZZ$ final states, dubbed $\langle\sigma v\rangle_{VV}$, or into $b\bar{b}$  [^4] There are currently no published equivalent exclusion lines for decays going mainly into $c\bar{c}$ nor $h_{1}h_{1}$.. For each point in parameter space, the procedure is to check the annihilation channel and apply the respective exclusion bounds in Fig. 1. We have confirmed that for the mass region studied, this constraint does not exclude a significant amount of the parameter space when applied after demanding the correct relic density.
Figure 1: Combined relevant limits from indirect searches on the total $\langle\sigma v\rangle$ as a function of the mass of the DM candidate $m_{DM}$. The lines coming from Fermi-LAT Ackermann and others (2015) and H.E.S.S. Abdalla and others (2022) assume a Navarro-Frenk-White (NFW) DM density profile and the AMS-02 Aguilar and others (2016) lines correspond to the conservative approach derived in Ref. Reinert and Winkler (2018).